OpenStudy (anonymous):

A store mixes Brazilian coffee worth $12 per kilogram and Turkish coffee worth $15 per kilogram. The mixture is to sell for $14 per kilogram. Find how much of each should be used to make a 444 kilogram mixture.

OpenStudy (anonymous):

put the number of kilos of turkish coffee as say \(x\) making the amount of brazilian coffee \(300-x\) since the total is 300

OpenStudy (anonymous):

the total dollars for the turkish coffee will be \(14x\) and the total dollars for the brazilian will be \(10(300-x)\) and you want that to amount to \(\$11\times 300=\$3300\) so you can set \[14x+10(300-x)=3300\] and solve for \(x\)

OpenStudy (anonymous):

what would ne my next step

OpenStudy (anonymous):

I dont get it.

OpenStudy (anonymous):

i got 75 for an answer but it is wrong

OpenStudy (anonymous):

the first step is to set up the equation if that is not clear, let me know where the confusion is the second step is solving the equation, and that is just mechanics, no thinking involved

OpenStudy (anonymous):

ok first off the question asks "Find how much of each " so your answer has to be two numbers, not one

OpenStudy (anonymous):

to solve \[14x+10(300-x)=3300\] first you have to remove the parentheses, by multiplying using the distributive law you get \[14x+3000-10x=3300\]

OpenStudy (anonymous):

the next step is to combine like terms on the left i.e. subtract \(14x-10x\) so you should get \[4x+3000=3300\]

OpenStudy (anonymous):

then subtract \(3000\) from both sides giving \[4x+=300\] and finally divide \[x=300\div4=75\] so yes, the answer is \(x=75\) meaning 75 pounds of turkish and the rest (225 pounds) of brazillian

OpenStudy (anonymous):

Oh ok that is where i messed up at.

OpenStudy (anonymous):

For what it's worth, I don't think this solution is correct. It is unclear where the number "300 kg" entered the discussion, but this doesn't come from the problem... the total of the two types of coffee is 444 kg, not 300. B + T = 444 <<-- total weight = # kg of Brazilian + # kg of Turkish and 12B + 15T = (14)(444) <-- per kg price of Brazilian X qty B plus price of Turkish X qty T equals price $14 X total weight 444 Sub in B = 444 - T: 12(444-T) + 15T = (14)(444) (12)444) - 12T + 15T = 14(444) 3T = 2(444) = 888 T = 296 <<-- and B = 444-T, so B = 148. In other words, 296 kg of Turkish, 148 kg of Brazilian. Check: weight: B+T = 148 + 296 = 444 kg total. and price: 12(148) + 15(296) = 6216 which is equal to 444kg x $14/kg

OpenStudy (anonymous):

wow that is weird. i am certain it was 300 at the start